OpenStudy (littlediva12459):

Will medal and fan plz help what is 12500/83 simplified

OpenStudy (xofallenangelox):

Radical? Or divided?

OpenStudy (xofallenangelox):

Are there any answer choices?

OpenStudy (xofallenangelox):

If it's divided, the answer is 150.6... and simplified would be 3.0

OpenStudy (littlediva12459):

The question is A park ranger tags 100 animals use a proportion to estimate the total population for the sample. 83 out of 125 animals are tagged So I did 83 over 125 equals 100 over x 83/125 100/x 83x=100x125 83x=12500 83x/83=125000/83 x=?

OpenStudy (littlediva12459):

The options are 151 animals 142 animals 67 animals 95 animals

OpenStudy (littlediva12459):

@xoFallenAngelox

OpenStudy (xofallenangelox):

Uhh o.o

OpenStudy (xofallenangelox):

x = 150.6

OpenStudy (xofallenangelox):

Just round the answer. If the number next to 6 is less than 5, then let it stay the same and say 150.6

OpenStudy (xofallenangelox):

And then 6 is over 5, so then round that to 151

OpenStudy (xofallenangelox):

So the amount is approximately 151 animals :)

OpenStudy (littlediva12459):

So the answer would round to 151 animals

OpenStudy (xofallenangelox):

Yes!

OpenStudy (littlediva12459):

thanks for your help i dont think we want .6 of an animal lol

OpenStudy (xofallenangelox):

LOL Sorry! I think that was my mistake XD You're welcome.

OpenStudy (littlediva12459):

well the question said to round so thanks

OpenStudy (xofallenangelox):

Yey >w< yw